Work out the price elasticity of demand for tickets from Lusaka to Livingston between 200 kwacha and 250


1
Expert's answer
2021-09-08T18:38:11-0400

"Q \n\nd\n\n\u200b\t\n\n =2000\u22125P"

"P=200 kwacha,"

"Q_d=2000-(5 \times 200)=1000"


"Old price quantity demanded=1000"


If "P=250 kwacha"

"=2000\u2212(5\u00d7250)=750"

"New price quantity demanded=1500""Price elasticity of demand =\frac{\frac{(Q_2-Q_1)}{(Q_1+Q_2)/2}}{\frac{(P_2-P_1)}{(P_2+P_1)/2}}""= \frac{\frac{(750-1000)}{(750+1000)/2}}{\frac{(250-200)}{(250+200)/2} } \\ =-1.3"